OpenStudy (anonymous):

use point slope formula to find the y-intercept. the points are (1,2) and (-2,4). i found the slope, it is -2/3.

OpenStudy (anonymous):

i started doing it but then when i came to 2y-4 = -4/3x - 4/3 i am stuck

OpenStudy (anonymous):

can you work it out

OpenStudy (ivettef365):

slope formula y2 - y1 ------- x2 - x1 you have points (1,2) and (-2,4) (x1,y1) (x2,y2)

OpenStudy (anonymous):

i already did that.. i got the slope.. it is -2/3.

OpenStudy (anonymous):

now i need to find the y-intercept so i can write the equation of the line in standard form.

OpenStudy (ivettef365):

ok, I see you got the slope

OpenStudy (ivettef365):

no you need to find the value of b on formula y = mx + b so you have m= -2/3 use any of the 2 points for x and y and find b

OpenStudy (ivettef365):

I meant now

OpenStudy (anonymous):

im using the point- slope formula..

OpenStudy (anonymous):

so u can use that instead?

OpenStudy (ivettef365):

oh, ok so you are using y - y1 = m(x - x1)

OpenStudy (anonymous):

no nevermind

OpenStudy (anonymous):

ill use the easier way

OpenStudy (anonymous):

ok , so i have 2= -2/3(1)+b

OpenStudy (ivettef365):

it is easier with y = mx + b

OpenStudy (ivettef365):

that is good so what is the value of b

OpenStudy (anonymous):

did you find b yet ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i was away

OpenStudy (anonymous):

how do you get it, i dont understand

OpenStudy (anonymous):

is -2/3 *1 -2/3 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so now idk what to do

OpenStudy (anonymous):

2= -2/3 +b

OpenStudy (anonymous):

y = mx + b 2 = -2/3(1) + b 2 = -2/3 + b (add 2/3 to each side) 2 + 2/3 = b 6/3 + 2/3 = b can you get it now ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

8/3 = b

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now how do i put it in standard form

OpenStudy (anonymous):

points are (1,2) and (-2,4)

OpenStudy (anonymous):

oh wait we dont need that info anymore

OpenStudy (anonymous):

slope is -2/3 and y-int is 8/3

OpenStudy (anonymous):

yes

OpenStudy (phi):

you now have y = -2/3 x + 8/3 One good first step is "clear the fractions" by multiplying both sides (and all terms ) by 3

OpenStudy (anonymous):

*update* guys, we've already found the y-intercept. now i need to write the standard form of the line.

OpenStudy (anonymous):

so, 3(y= -2/3+8/3) ?

OpenStudy (phi):

yes

OpenStudy (phi):

but there is an x in there next to the -2/3

OpenStudy (anonymous):

3y = -2x+8

OpenStudy (anonymous):

now divide by 3?

OpenStudy (anonymous):

Ax + By = C y = -2/3x + 8/3 (multiply by 3) 3y = -2x + 8 now add 2x to both sides and what do you get ?

OpenStudy (phi):

standard form is Ax + By = C A is positive, and an integer. no, dividing by 3 is "going backwards" just add 2x to both sides and simplify

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so i add 2x to 3y and 8?

OpenStudy (anonymous):

-2x i mean

OpenStudy (phi):

like this 2x+ 3y = -2x + 8 +2x

OpenStudy (anonymous):

now what

OpenStudy (anonymous):

you see how the 2x's cancel each other out on the right side

OpenStudy (phi):

on the right side you have 2 x's take away 2 x's that can be simplified to 0

OpenStudy (anonymous):

now what are you left with

OpenStudy (anonymous):

2x+3y = 8

OpenStudy (anonymous):

you got it :)

OpenStudy (anonymous):

finally :)

OpenStudy (anonymous):

good job

OpenStudy (phi):

Just for reference, you said ***when i came to 2y-4 = -4/3x - 4/3 i am stuck*** First, that should be \[y- 4= -\frac{2}{3}\left( x - (-2)\right) \\ y- 4= -\frac{2}{3}\left( x +2\right) \\ y-4 = -\frac{2}{3}x -\frac{4}{3}\] if you are trying to put that into y = mx+b form, you want the y "by itself" the first step is add +4 to both sides \[ y-4+4= - \frac{2}{3}x - \frac{4}{3} +4 \] the left side simplifies to y the right side has (unfortunately) fractions. Make 4 an improper fraction by multiplying by 3/3 you get \[y= - \frac{2}{3}x - \frac{4}{3} + \frac{12}{3} \] combine 12/3 - 4/3 to get 8/3: \[ y= - \frac{2}{3}x + \frac{8}{3} \]